Present and future value tables of $1 at 3% are presented below:

N

FV $1

PV $1

1

1.03000

0.97087

2

1.06090

0.94260

3

1.09273

0.91514

4

1.12551

0.88849

5

1.15927

0.86261

6

1.19405

0.83748

7

1.22987

0.81309

8

1.26677

0.78941

9

1.30477

0.76642

10

1.34392

0.74409

11

1.38423

0.72242

12

1.42576

0.70138

13

1.46853

0.68095

14

1.51259

0.66112

15

1.55797

0.64186

16

1.60471

0.62317

Joseph wants to invest in a 6% CD account that compounds semiannually. He would like this account to have a balance of $100,000 four years from now. How much must Joseph deposit to accomplish his goal?

$88,849

$78,941

$25,336

$22,510
Answer #1

Option B

78,941

Amount to be deposited

= 100,000 * PV $1 @3%, n= 8

= 100,000 * 0.78941

= 78,941